OpenStudy (jpes2193):

I'm working on fractional eponents... does any one have some simple tips or tricks to help? Im really struggling with the concept.

OpenStudy (misty1212):

\[\huge \sqrt[n]{x^m}=x^{\frac{m}{n}}\]

OpenStudy (nincompoop):

fractional exponents are another way of writing radicals

OpenStudy (jpes2193):

so what do I do with \[x ^{3/10}*x ^{1/5}\] I know that we add the exponets so i would have to take those and find the LCD, then add them, then do the rewritting of the radical?

OpenStudy (agent0smith):

Yes, exactly.

OpenStudy (nincompoop):

\(\huge x^\frac{p}{q} \times x^\frac{a}{b} \rightarrow x^{\frac{p}{q}+\frac{a}{b}} \)

OpenStudy (nincompoop):

you don't necessarily have to write it into radical form unless you're asked to do so

OpenStudy (jpes2193):

so my first step with this would be to distribute \[(-5x ^{3/4}y)(-2x ^{1/2}y ^{2/3})\]

OpenStudy (jpes2193):

The teacher says that we should write the answer in what ever form the question is being asked in...

OpenStudy (nincompoop):

distribute what?

OpenStudy (jpes2193):

the group on the left to the group on the right

OpenStudy (nincompoop):

one trick is to make all the exponents of common variables with the same denominator so \(\large y^\frac{3}{3} \times y^\frac{2}{3} \) since in fractional form \(\frac{p}{p}=1\)

OpenStudy (nincompoop):

\(\large \frac{3}{4} + \frac{1}{2} \rightarrow \frac{3}{4} + \frac{2}{\color{red}{4}} \)

OpenStudy (jpes2193):

so \[x ^{3/4}*x ^{2/4}\]

OpenStudy (nincompoop):

correct

OpenStudy (jpes2193):

so it would be \[x ^{5/4}\]

OpenStudy (nincompoop):

see how it easier it is?

OpenStudy (nincompoop):

now don't forget the coefficients, because they play a huge part specially the signs

OpenStudy (jpes2193):

yes...

OpenStudy (jpes2193):

\[10x ^{5/4}y ^{5/3}\] would be the answer

OpenStudy (nincompoop):

you know that what I did is just the same as looking for the LCD, I just reworded it in a way it fits right in your problem example

OpenStudy (nincompoop):

correct

OpenStudy (jpes2193):

thank you so much for your help :) fractions are always a pain.. haha

OpenStudy (nincompoop):

ye they are, but they are very important

OpenStudy (nincompoop):

you can train yourself to solve for fractions fast by doing a lot of problems, first by hand and then force yourself to do it mentally